acopeland27:

Solve a/-6 + 8 = 12.

MusicGeek:

So, I don't know how to do this. And im bein honest bout that. but when i googled it, this is what it gave me: a/(-6)+8 = 12 // - 12 a/(-6)-12+8 = 0 -1/6*a-4 = 0 // + 4 -1/6*a = 4 // : -1/6 a = 4/(-1/6) a = -24 a = -24

acopeland27:

that works

acopeland27:

=)